2. (5 points) Nerwin, Inc. is a furniture manufacturing company with 50 employees. Recently, after a long negotiation with th2. (5 points) Nerwin, Inc. is a furniture manufacturing company with 50 employees. Recently, after a long negotiation with the local labor union, the company decided to initiate a pension plan as a part of its compensation plan. The plan will start on January 1, 2014. Each employee covered by the plan is entitled to a pension payment each year after retirement. On the basis of a discussion with the supervisor of the Personnel Department and an actuary from an insurance company, the controller develops the following information related to the pension plan.Average length of time to retirement10 yearsExpected life duration after retirement15 yearsTotal pension payment expected each year after retirement for all employees. Payment made at the end of the year. Amount includes all employees.$300,000 per yearThe interest rate to be used is 7%.Required:Calculate the present value of this deferred annuity (at January 1, 2014).
Answer #1

Answer :-
Calculation of Present value of annual pension payment at the end of 15th year :-

= 300000* [ 1 - (1 + 0.07)-15 ] / 0.07

= 300000* [ 1 - (1.07)-15 ] / 0.07

= 300000* [ 1 - 0.36245] / 0.07

= 300000* 0.63755 / 0.07

= 191265 / 0.07

= $ 2732357.14 (approx)

Calculation of Present value of deferred annuity at the starting of pension plan i.e., on January 1, 2014

= 2732357.14 / (1.07)10

= 2732357 / 1.9672

= $ 1388957 (approx)

Conclusion :- Present value of deferred annuity on January 1, 2014 = $ 1388957 (approx).